UT September 09, 2019 - Restart Night #8

OSA: S Allanson
Observer: O Kuhn

Summary:

  • MOS acquisition to check RZP. The old one is fine for L1 and the new one determined a few days ago, 177.30 deg, for L2. (Reverted to old value at the end of the night. PCS will need to be restarted to take it).
  • Seeing was too poor to collect transform data.
  • LUCI NS guiding + dithering test, to check GCS patch 20190908 to fix IT 7845. We did not see any "guide loop did not pause within timeout" errors. However, the shorter WFS exptime that was manually set via the Guide Control gui did not stick. It was increased dynamically when counts were deemed too low (as they were because of clouds + poor seeing). Note added to IT 7643.
  • Collected flexcomp data for G200+N1.8, G210+N3.75 and G200+N3.75. Dave installed flexcomp tables for Imaging and G210+N1.8. Collected data to verify flexcomp table for G210+N1.8.

Log:

01:38 - Before sunset, it was pretty clear, much better than anticipated. But at sunset, some low clouds are billowing over. Waiting for these to pass before opening.

Initialized both LUCI1 and LUCI2. Running these on obs3, and bringing up RTDs using script ./startRTD.sh [1|2].

02:11 - Looking better but still some virga to the south. Putting the Astrometric M52 mask in, in prep for rotator zeropoint checks. Will update PCSInstrument.conf to use RIGHTZEROPOINT 177.30 before the check.

02:15 - LUCIFER RIGHT ROTATORZEROPOINT set to 177.30 in PCSInstrument.conf
cp PCSInstrument.20190907_NEW_LUCI_RZP.conf PCSInstrument.conf
Steve restarted PCS

02:16 - Opening the enclosure. While moving the mask to the astrometric mask, there was a LUCI2 MOS error: Error while executing seq: MHU picker arms - grab mask. Grabber error. Called Dave and he will recover it but is ~10 min from home. He has given us the go ahead to point and collimate, meanwhile.

02:40 - Seeing is terrible, at least at the elevation of M52 (35 deg). It is 2.5" !! Like this, no way to collect transform data. We can run the NS test and later, 2nd half, we might go on to LUCI spec tasks.

02:50 - sending M52 MOS acq script. sky/obj are 331/332, mask 333. Results:
  • L2: sw had mis-id'd 34. id'd it manually, then dx,dy,dtheta = 2.55252, -3.20889, 0.02004 (error 0.0635,0.08963). Mask-to-det: X-shift,Y-shift,angle = 27.8, -4.66, -0.2478.
  • L1: sw had mis-id'd 34 (box) and 35 (no box). id'd both manually, then: dx,dy,dtheta = 4.63649, -1.46804, 0.09686 (error 0.07828, 0.09486). Mask-to-det: X-shift, Y-shift, angle = -2.0256, 0.6050, 0.0298.

Clicked "Send" for for L2 after I finished the alignment, but got an error: "Could not send data from RTD to Scheduler." Running LUCI on obs3 and RTD via ./startRTD.sh 2 in /lbt/lucifer/lcsp/bin/ Same for L1. ?? Sending offsets from the telsvc gui. (This is known --- when the RTD GUIs are not run up by clicking the button on the observer panel, they do not communicate with the sequencer, so they cannot send offsets.)

As Jenny noted before, the overlays depicting the box centers do not fall exactly on the box center after the mask-to-detector alignment.

03:13 Taking confirmatory image, 334. Seeing better, but still poor: 1.5"-1.8". On L1, there is a hint that the stars are below Y-centers of the boxes at the top - see the ones on the right and left corners, but and slightly above center in the boxes at bottom. On L2, it is similar. I am looking at the difference 334 minus 332.

03:27 Based on the pair of acquistions, not changing zeropoints - just confirming they are OK. Slewing to Stone_L1 to check seeing here. It was still out-of-spec for transform data collection at M52, but that was low.

03:35 Slewing to Stone_L. It's pretty close to the moon. Seeing is bouncing around still - from 1 to 2". Going to do the NS guiding + dithering... to test patch GCS 20190908 (fix IT 7845).

03:40 NS guiding + dithering test using 2014 AF51. Used twin.py to convert L2 script (works fine, but maybe filter tables need to be updated in the script?).

03:50 Slew to field of 2014 AF51. Forgot to "Set". Put pause in. Also forgot to turn off flex comp lookup table during slew.

03:54 Resending preset - non-sidereal, this time. Checking the patch Petr installed. We set exp time to 4 sec via Guide Control. WFS is OK on the left side, but on the right side, we don't see spots - WFS "spots too dim or image saturated". Same star (guide flux is the same). Increased WFS exptime to 18 on DX, but still no spots.

04:01 Clicked ABORT on the observer panel while both L1 and L2 were integrating. Sent abort to L1 but not to L2. Clicked ABORT again. This time readout on L2 was aborted.

04:04 Because of the DX WFS images not showing spots, we are restarting AGW-Cam2 computer and GCS. Going to pointing star to confirm that this worked. It did. We see spots on both sides.

04:10 Problem with WFS turbo mode taking over. We are trying to set exptime to 5-sec, but watching as the WFS exptimes ramp up to 18 and 30 sec. Getting the red BLOCKED messages. Clicking set each time it ramps up is not consistent with the fix reported in IT , is it? Steve is going to tweak turbo mode.

04:12 Clouds moved in. Lost the star due to clouds.

04:15 Steve tweaked turbo in GCS.conf to made it 5-sec. This is too short for this star+clouds. Now he's trying 9-sec. This is related to IT 7643 and IT 6540 (cannot set WFS exptime in GCS GUI).

04:22 Summary: no guider timeouts (IT 7845 fixed), but we have problems with setting WFS exptime (IT 7643 - Steve added a note to it.)

04:35 Closed for threatening cloud. RH also spiking. 88% now. Flex Comp data at G200 + N1.8 needs to be repeated. We may start that in a bit, with the caveat that we may need to repeat if we open before it's complete.

04:47 I had clicked "LUCI" on rm507-5 to run up LUCI a while ago, after the astrom mask acquisition. That brought up the RTD GUIs. Since that time, no new images have been displayed to the Aladins that I opened earlier from the RTDs on obs3.

05:00 Opened LUCI Observer and all java GUIs on 507-5. Copied the flex comp scripts to rm507-5:/home/observer/share (this should be going to telescope account, though...)

05:16 Configuring for Flex Comp N1.8+G200:
  • Moving in OptSpecSieve (MOS error on L2 - MHU picker arms - grab mask...) Rotator was at 67 deg. Dave fixed the problem and now, with rotator at 210, it grabbed mask. But, it did an auto-correct.

05:38 Dave finished reducing the G210+N180 flex comp data and will update that table, so that we can do some spec checkout in case we reopen. Waiting for that before starting the flex comp data collection.

NB: before starting flex comp data collection, set both rotators to -89 deg. If they are at park, they may take too long to reach -89 and the script will timeout.

05:57 Dave updated the flex comp tables in the LUCI database for N375+MIRR & G210+N180.

06:00 Steve moved rotators to -89 deg. Dome closed & dark. Rested the shells, since we need to go to 24 deg elev.

06:00 Taking a dark with blind (note that when checking headers, filter1=blind but filter2=J). L1 338, L2 338

06:00 Deploying cal unit.

06:09 At zenith. Starting FlexComp_Calib15.xml script.

Spots look pretty aberrated on L1, better on L2.

Init detector focus and toggle filter from zJspec to J and back to zJspec.

06:15 Start over. Script is sending rot from 15 back to -90. First exposure 347.

06:17 Restart, just in case rotator did not settle. First exposure 348. Same aberration with L1. Going to continue as the aberration is field dependent, meaning it's better in some regions, and the relative positions of spots can be determined still. Compared an image from tonight with one from this afternoon - they both have similar aberrations.

07:02 zen complete: 348-420

07:02 68 deg elev: 421-493

07:44 46 deg elev: 494-566

08:22 24 deg elev: 567-639

08:59 a dark, 640, this time with 10 ndits (1 x 10 x 5-sec, LIR/INT)

FLEX COMP VERIFICATION FOR G210 + N1.8, J, gratwlen = 1.25

Config:
  • N1.8
  • OptSpec Sieve
  • G210 @ 1.25 mic
  • clear + J
  • flexcomp ON (LOOKUP)
  • argon ON
  • FlexComp_Verify54d.xml

09:06 35 deg elev: Running FlexComp_Verify54d.xml
  • FC Verify 35 deg: 641-662
  • note that aberrations on L1 are gone now, with this config.

~09:30 It has started to rain.

09:31 Turn off FlexComp on Instrument GUI and slew to 79 deg elev, turn FlexComp back on and run FlexComp_Verify54d.
  • FC Verify 79 deg: 663-682

FLEX COMP FOR G210 + N3.75, J, gratwlen = 1.25

09:57 Setup for G210+N3.75: Turn off FC and slew to 90 deg.
  • N3.75
  • OptSpec Sieve
  • G210 @ 1.25 mic
  • clear + J
  • flexcomp OFF
  • argon ON
  • FlexComp_Calib15.xml (takes 5.0sec x 1 x 1)

10:02 Zen 683-755
  • Spots look a bit elongated in x direction on L1, but not L2.

10:37 elev 68 deg: 756-828

11:15 elev 46 deg: 829-901

11:52 elev 24 deg: 902-974
  • L1 taking data while NOT STABLE on observer panel (probably has happened earlier too)

FLEX COMP COLLECTION FOR G200 + N3.75, zJspec, gratwlen = 1.17
  • N3.75
  • OptSpec Sieve
  • G200 @ 1.17 mic
  • clear + zJspec
  • flexcomp OFF
  • argon ON
  • FlexComp_Calib15.xml (takes 5.0sec x 1 x 1)

12:35 24 deg elev: 975-1047
  • L1 still looks elongated horizontally, L2 ok

13:09 46 deg elev: 1048-1120
  • sometimes L1 integrating while NOT STABLE

13:21: Reverted to old L2 RZP in PCSInstrument.conf. This was not done for the LUCI observing after the RZP confirmations, that is for the NS tests, but the difference between old and new L2 RZPs is small, 0.25 deg, and should not have affected the NS testing.

13:50 68 deg elev: 1121-1193

14:25 90 deg elev: 1194-

Plan Outline:

Plan rolled over from Night #7
  • LUCI
    • AGW-1 Transform
      • No need to update PCS, rotator ZP looks good after astrometric mask test
    • IF spectroscopic modes are released for LUCIs
      • G200 science-like OB
      • G210 science-like OB
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
  • MODS
    • Test NS guiding with dithering

Closed-Dome Tasks:
  • Collection of LUCI1+2 Flex Comp Data
    • N1.8+G200 and N1.8+G210 were done this afternoon. N3.75 Imaging was done earlier.
    • Dave wrote: "use the script in /home/dthompson/LUCI/Flexure_Compensation called FlexComp_Calib15.xml. There is a README file there as well that describes more how to take the data. Basically, you set up the instruments manually as needed then run the script (it just sets the rotators and takes images)."
    • Remaining TBD:
      • N375+G210
      • N375+G200
      • N30 + MIRR (CU/Halo1 for enough counts?)
      • N180 +MIRR
      • Lower priority for now (G040 not released, G150 seldom used)
        • N375 + G150/G040
        • N180 + G150/G040
        • N30 + G150/G040
      • Eventually:
        • N30 + G210 - low priority (won’t be used much)
        • X N30 + G200 # This mode is not used, grating not AO quality
  • Check MODS offsets from GUI:
    • with MODS authorized, to confirm that, to send a relative offset from the MODS GUI after the script has executed an absolute offset, one must reselect REL on the GUI. Likewise with the coordinate system - DETXY vs RADEC.
The GUI always shows DETXY and REL, but if the script has sent something else, then it is necessary to reselect DETXY and/or REL if you want DETXY and/or REL.
  • Continue taking MODS calibrations.
  • Check MODS focus. Already spectra and images were obtained in all modes and focus looks OK, except for the one of the prism modes. (Examination of standards taken in all spec modes on 20190904 shows no strong evidence for defocus. Seeing fluctuations bounced the spatial profile around, from 7 - 11 pix. We noticed the 11 in a prism spectrum, but it was due to fluctuations.
-- BarryRothberg - 08 Sep 2019
Topic revision: r6 - 09 Sep 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback